# The Curse of Depth in Large Language Models

### v2

@@ 5.2 Supervised Fine-tuning
To verify whether the gains in pre-training can be translated to the stage of post-training, we perform SFT with the models obtained from Section 5.1 on the Commonsense170K dataset Hu et al. (2023) across eight downstream tasks. We adopt the same fine-tuning configurations as used in Li et al. (2024b). The results, presented in Table 2, demonstrate that LayerNorm Scaling consistently surpasses other normalization techniques in all evaluated datasets. For the LLaMA-250M model, LayerNorm Scaling improves average performance by 1.80% and achieves a 3.56% gain on ARC-e compared to Mix-LN. Similar trends are observed with the LLaMA-1B model, where LayerNorm Scaling outperforms Pre-LN, Post-LN, Mix-LN, and DeepNorm on seven out of eight tasks, with an average gain of 1.86% over the best baseline. These results confirm that LayerNorm Scaling enhances generalization on diverse downstream tasks by improving the representation quality of deep layers.
@@
Post-LN (Ba, 2016) 22.95 37.83 26.94 52.72 26.17 11.60 49.56 32.54
DeepNorm (Wang et al., 2024) 23.60 37.86 36.62 61.10 25.69 15.00 49.57 35.63
Mix-LN (Li et al., 2024b) 26.53 56.12 41.68 66.34 30.16 18.00 50.56 41.34
Pre-LN (Baevski and Auli, 2019) 24.93 38.35 40.15 63.55 26.34 16.20 49.01 36.93
@@
Post-LN (Ba, 2016) 22.95 37.82 25.08 49.51 25.04 13.80 49.57 31.96
DeepNorm (Wang et al., 2024) 23.35 37.83 27.06 52.94 26.19 11.80 49.49 32.67
Mix-LN (Li et al., 2024b) 23.19 37.83 25.08 49.51 25.04 11.80 49.57 31.72
Pre-LN (Baevski and Auli, 2019) 26.54 62.20 45.70 67.79 30.96 17.40 50.51 43.01
@@ 5.3.1 OLMo
Model Size Scaling. To further assess the scalability and robustness of LNS, we conduct additional experiments using the OLMo repository (Groeneveld et al., 2024), scaling training across model sizes of 60M, 150M, 300M, 1B, and 7B parameters. All models are trained on a fixed 20B-token budget to ensure comparability. These experiments are designed to evaluate whether the performance gains observed with LNS in smaller-scale settings extend to more challenging and state-of-the-art LLM training regimes. As shown in Figure 1, LNS consistently and substantially outperforms the standard Pre-LN baseline across all model sizes. Remarkably, for the 7B model, LNS reduces the final loss from 2.69 to 2.50. These results underscore the scalability of LNS and its effectiveness in large-scale pre-training scenarios.
@@
Beating OLMo’s Scaled Initialization. OLMo adopts the scaled initialization proposed in Zhang et al. (2019) and used by Mehta et al. (2024), which scales input projections by $1/\sqrt{d_{\mathrm{model}}}$, and output projections by $1/\sqrt{2\cdot d_{\mathrm{model}}\cdot l}$ at every layer. This method is designed to enhance training stability and to scale down variance at initialization. To evaluate the effectiveness of LNS, we compare it against this state-of-the-art initialization by training OLMo-1B on 20B tokens. As shown in Table 3, LNS achieves consistently lower training loss, indicating that it may offer a more effective alternative for large-scale LLM training.
@@
We further evaluate the generalizability of LNS by applying it to a state-of-the-art architecture, Qwen2.5-0.5B (Yang et al., 2024). We train the model for 6B tokens and compare LNS against the standard Pre-LN setup. Consistent with previous findings, Table 4 illustrates that LNS yields a notable reduction in perplexity—from 20.62 to 19.57—highlighting its effectiveness even on strong, modern architectures.
@@ 5.4 LNS Effectively Scales Down Output Variance
As LNS is proposed to reduce output variance, we empirically validate this claim during the pre-training of LLMs. We compare the layerwise output variance of three configurations: (1) the standard Pre-LN (Ba, 2016), (2) Pre-LN with Scaled Initialization (Shoeybi et al., 2020; Radford et al., 2019), which scales the initialization of the feedforward layers’ weights $W_{0}$ and $W_{2}$ by $\frac{1}{\sqrt{2L}}$, where $L$ is the total number of Transformer layers, and (3) Pre-LN with LNS. The average output variance across layers is shown in Figure 4. For both vanilla Pre-LN and Scaled Initialization, the output variance in shallow layers (blue) remains relatively stable throughout training, while variance in deeper layers (red) grows substantially after 2K iterations, reaching up to 175 in the final layer. Since Scaled Initialization only operates at initialization, it is insufficient to constrain output variance during training. In contrast, LNS consistently suppresses the growth of output variance in deeper layers, capping it at approximately 25.
@@ 5.6 LayerNorm Scaling in Vision Transformer
To evaluate whether LNS also benefits architectures beyond language models, we conduct experiments on ViT-S on ImageNet-1K. Since ViT-S includes LayerScale (Touvron et al., 2021) by default—which may interfere with the effect of LNS—we remove LayerScale from all evaluated variants to ensure a fair comparison. We then test different insertion positions of LNS. The top-1 accuracy results are summarized in Table 5. Whereas LNS in language models is typically most effective directly after normalization, in Vision Transformers, the best position is after the attention and MLP blocks. We next examine whether this performance gain correlates with better control of layer-wise variance.
@@ 6 Ablation Study
Comparing Against Other Scaling Methods. We first compare LNS with previous scaling approaches, including (1) Scaled Initialization (Shoeybi et al., 2020; Radford et al., 2019), which scales the initialization of $W_{0}$ and $W_{2}$ by the overall depth $1/\sqrt{2L}$; (2) Depth-Scaled Initialization (Zhang et al., 2019) scales the initialization of weight matrices by the current depth $1/\sqrt{2l}$; (3) SkipInit (De and Smith, 2020) introduces a learnable parameter after FFN/Att layers, initialized as $1/\sqrt{L}$; (4) LayerScale (Touvron et al., 2021) applies per-channel weighting using a diagonal matrix, diag$(\lambda_{1},\ldots,\lambda_{d})$, where each weight $\lambda_{i}$ is initialized to a small value (e.g., $\lambda_{i}=\epsilon$). Table 6 presents the results of LLaMA-130M and LLaMA-250M.
@@
+ LayerScale (Touvron et al., 2021) 27.93 23.45
+ SkipInit (De and Smith, 2020) 27.41 22.29
+ Depth-Scaled Initialization (Zhang et al., 2019) 26.95 21.50
+ Scaled Initialization (Shoeybi et al., 2020) 26.04 20.98
@@
Comparison with Other Layer Normalization. In addition, we conducted comparisons using LLaMA-130M to evaluate LayerNorm Scaling against recently proposed normalization methods, including Admin (Liu et al., 2020), Sandwich-LN (Ding et al., 2021), Group-LN (Wu and He, 2018; Ma et al., 2024), and Mix-LN (Li et al., 2024b). Table 7 shows that Admin and Group-LN degrade performance. Sandwich-LN slightly outperforms Pre-LN. Both Mix-LN and LayerNorm Scaling improve over Pre-LN by good margins. However, Mix-LN fails to reduce perplexity under 26, falling short of LayerNorm Scaling and suffers from instability in large-scale scenarios as shown in Table 1.
@@ 7 Related Work
Ineffectiveness of Deeper Layers in Transformers. The ineffectiveness of deep layers in LLMs has been previously reported. Yin et al. (2024) found that deeper layers of LLMs can tolerate significantly higher levels of pruning compared to shallower layers, achieving high sparsity. Similarly, Gromov et al. (2024) and Men et al. (2024) demonstrated that removing early layers causes a dramatic decline in model performance, whereas removing deep layers does not. Lad et al. (2024) showed that the middle and deep layers of GPT-2 and Pythia exhibit remarkable robustness to perturbations such as layer swapping and layer dropping. Recently, Li et al. (2024a) highlighted that early layers contain more outliers and are therefore more critical for fine-tuning. While these studies effectively highlight the limitations of deep layers in LLMs, they stop short of identifying the root cause of this issue or proposing viable solutions to address it.
@@
Layer Normalization in Language Models. LN (Ba, 2016) was initially applied after the residual connection in the original Transformer (Vaswani, 2017), which is known as Post-LN. Later on, Pre-LN (Baevski and Auli, 2019; Dai et al., 2019; Nguyen and Salazar, 2019) dominated LLMs, due to its compelling performance and stability (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023; Bi et al., 2024). Prior works have studied the effect of Pre-LN and Post-LN. Xiong et al. (2020) proves that Post-LN tends to have larger gradients near the output layer, which necessitates smaller learning rates to stabilize training, whereas Pre-LN scales down gradients with the depth of the model, working better for deep Transformers. Wang et al. (2019) empirically confirmed that Pre-LN facilitates stacking more layers and Post-LN suffers from gradient vanishing. The idea of connecting multiple layers was proposed in previous works (Bapna et al., 2018; Dou et al., 2018; Wang et al., 2019). Admin introduces additional parameters to control residual dependencies, stabilizing Post-LN. DeepNorm (Wang et al., 2024) enables stacking 1000-layer Transformers by upscaling the residual connection before applying LN. Additionally, Ding et al. (2021) proposed Sandwich LayerNorm, normalizing both the input and output of each transformer sub-layer. Takase et al. (2023a) introduced B2T to bypass all LN except the final one in each layer. Li et al. (2024b) recently combines Post-LN and Pre-LN to enhance the middle layers. Zhu et al. (2025b) introduces Dynamic Tanh (DyT) as a normalization-free alternative in Transformers, delivering comparable performance. Zhuo et al. (2025) proposes HybridNorm, a hybrid normalization scheme combining QKV normalization with Post-Norm FFN to stabilize training in deep transformers. De and Smith (2020) also states that normalized residual blocks in deep networks are close to the identity function and proposes SkipInit to remove normalization by introducing a learnable scalar multiplier on the residual branch initialized to $1/\sqrt{L}$. Our experiments suggest that SkipInit’s learnable parameter does not improve performance and sometimes harms training.
@@ Appendix A Proofs of the Theorems of curse of depth
Given Equation (2) from Takase et al. (2023b), we have:
@@ A.1.2 Variance of the Feed-Forward Network
Using the result obtained by Wang et al. (2024), we get:
@@
This provides a tighter lower bound for ${\sigma^{2}_{x_{\ell}}}$ compared to the upper bound of Equation (33). Since we know the upper bound of variance grows exponentially, the lower bound must be sub-exponential. Therefore, for ${\sigma^{2}_{x_{\ell}}}=\ell$, we must have:
@@
∎ Therefore, the increasing lower bound for ${\sigma^{2}_{x_{\ell}}}$ must grow faster than a linear function. So, the increase of variance is sub-exponential. A large increase in such bound will lead to gradient spikes, which can connect to previous studies in Huang et al. (2025a, b).
@@
×$\times\Big{(}1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\Big{(}\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big{)}\Big{)}.$
@@ A.2.1 Proof of Lemma 36
Our derivation follows results in Takase et al. (2023b), specifically Equation (7), which provides an upper bound on the norm of $\frac{\partial y_{\ell}}{\partial x_{1}}$ as:
@@
[…]
@@
Let $\sigma_{w1_{\ell}}$ and $\sigma_{w2_{\ell}}$ be the standard deviations of $W^{1}_{\ell}$ and $W^{2}_{\ell}$, respectively. From Assumption 3.1, the spectral norms of $W^{1}_{\ell}$ and $W^{2}_{\ell}$ are given by their standard deviations and dimensions (Vershynin, 2018), so wo have:
@@
From Vershynin (2018), we know that:
@@
×$\times\Big{(}1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\Big{(}\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big{)}\Big{)}.$
@@ A.2.2 Analysis of the Upper Bound
As discussed in Takase et al. (2023b), $\sigma$ should be sufficiently small, and the standard deviation, $\sigma_{x^{\prime}_{\ell}}$ or $\sigma_{x_{\ell}}$ should satisfy the condition $\sigma^{2}\ll\sigma_{x^{\prime}_{\ell}}$ to maintain the lazy training scheme. Thus, we obtain the following bound for the product over $\ell$ from 1 to $L$:
@@
Equation (36) can be expressed in a traditional product form Whittaker and Watson (1996) for $\sigma_{x_{\ell}}$:
@@
To establish a lower bound for ${\sigma^{2}_{x_{\ell}}}$, we exponentiate both sides. Setting ${\sigma^{2}_{x_{\ell}}}=\ell$, we must have:
@@
Here $\epsilon$ is a small constant with $11/2\leq\epsilon<$ 1. Therefore, the increasing upper bound of ${\sigma^{2}_{x_{\ell}}}$ is slower than the $\ell^{3}$ function, leading to:

### v3

@@ 5.2 Supervised Fine-tuning
To verify whether the gains in pre-training can be translated to the stage of post-training, we perform SFT with the models obtained from Section 5.1 on the Commonsense170K dataset hu2023llm across eight downstream tasks. We adopt the same fine-tuning configurations as used in li2024mix. The results, presented in Table 2, demonstrate that LayerNorm Scaling consistently surpasses other normalization techniques in all evaluated datasets. For the LLaMA-250M model, LayerNorm Scaling improves average performance by 1.80% and achieves a 3.56% gain on ARC-e compared to Mix-LN. Similar trends are observed with the LLaMA-1B model, where LayerNorm Scaling outperforms Pre-LN, Post-LN, Mix-LN, and DeepNorm on seven out of eight tasks, with an average gain of 1.86% over the best baseline. These results confirm that LayerNorm Scaling enhances generalization on diverse downstream tasks by improving the representation quality of deep layers.
@@
Post-LN (ba2016layer) 22.95 37.83 26.94 52.72 26.17 11.60 49.56 32.54
DeepNorm (wang2024deepnet) 23.60 37.86 36.62 61.10 25.69 15.00 49.57 35.63
Mix-LN (li2024mix) 26.53 56.12 41.68 66.34 30.16 18.00 50.56 41.34
Pre-LN (baevski2018adaptive) 24.93 38.35 40.15 63.55 26.34 16.20 49.01 36.93
@@
Post-LN (ba2016layer) 22.95 37.82 25.08 49.51 25.04 13.80 49.57 31.96
DeepNorm (wang2024deepnet) 23.35 37.83 27.06 52.94 26.19 11.80 49.49 32.67
Mix-LN (li2024mix) 23.19 37.83 25.08 49.51 25.04 11.80 49.57 31.72
Pre-LN (baevski2018adaptive) 26.54 62.20 45.70 67.79 30.96 17.40 50.51 43.01
@@ 5.3.1 OLMo
Model Size Scaling. To further assess the scalability and robustness of LNS, we conduct additional experiments using the OLMo repository (groeneveld2024olmo), scaling training across model sizes of 60M, 150M, 300M, 1B, and 7B parameters. All models are trained on a fixed 20B-token budget to ensure comparability. These experiments are designed to evaluate whether the performance gains observed with LNS in smaller-scale settings extend to more challenging and state-of-the-art LLM training regimes. As shown in Figure 1, LNS consistently and substantially outperforms the standard Pre-LN baseline across all model sizes. Remarkably, for the 7B model, LNS reduces the final loss from 2.69 to 2.50. These results underscore the scalability of LNS and its effectiveness in large-scale pre-training scenarios.
@@
Beating OLMo’s Scaled Initialization. OLMo adopts the scaled initialization proposed in zhang2019improving and used by mehta2024openelm, which scales input projections by $1/\sqrt{d_{\mathrm{model}}}$, and output projections by $1/\sqrt{2\cdot d_{\mathrm{model}}\cdot l}$ at every layer. This method is designed to enhance training stability and to scale down variance at initialization. To evaluate the effectiveness of LNS, we compare it against this state-of-the-art initialization by training OLMo-1B on 20B tokens. As shown in Table 3, LNS achieves consistently lower training loss, indicating that it may offer a more effective alternative for large-scale LLM training.
@@
We further evaluate the generalizability of LNS by applying it to a state-of-the-art architecture, Qwen2.5-0.5B (qwen2.5). We train the model for 6B tokens and compare LNS against the standard Pre-LN setup. Consistent with previous findings, Table 4 illustrates that LNS yields a notable reduction in perplexity—from 20.62 to 19.57—highlighting its effectiveness even on strong, modern architectures.
@@ 5.4 LNS Effectively Scales Down Output Variance
As LNS is proposed to reduce output variance, we empirically validate this claim during the pre-training of LLMs. We compare the layerwise output variance of three configurations: (1) the standard Pre-LN (ba2016layer), (2) Pre-LN with Scaled Initialization (shoeybi2019megatron; radford2019language), which scales the initialization of the feedforward layers’ weights $W_{0}$ and $W_{2}$ by $\frac{1}{\sqrt{2L}}$, where $L$ is the total number of Transformer layers, and (3) Pre-LN with LNS. The average output variance across layers is shown in Figure 4. For both vanilla Pre-LN and Scaled Initialization, the output variance in shallow layers (blue) remains relatively stable throughout training, while variance in deeper layers (red) grows substantially after 2K iterations, reaching up to 175 in the final layer. Since Scaled Initialization only operates at initialization, it is insufficient to constrain output variance during training. In contrast, LNS consistently suppresses the growth of output variance in deeper layers, capping it at approximately 25.
@@ 5.6 LayerNorm Scaling in Vision Transformer
To evaluate whether LNS also benefits architectures beyond language models, we conduct experiments on ViT-S on ImageNet-1K. Since ViT-S includes LayerScale (touvron2021going) by default—which may interfere with the effect of LNS—we remove LayerScale from all evaluated variants to ensure a fair comparison. We then test different insertion positions of LNS. The top-1 accuracy results are summarized in Table 5. Whereas LNS in language models is typically most effective directly after normalization, in Vision Transformers, the best position is after the attention and MLP blocks. We next examine whether this performance gain correlates with better control of layer-wise variance.
@@ 6 Ablation Study
[…]
@@
+ LayerScale (touvron2021going) 27.93 23.45
+ SkipInit (de2020batchnormalizationbiasesresidual) 27.41 22.29
+ Depth-Scaled Initialization (zhang2019improving) 26.95 21.50
+ Scaled Initialization (shoeybi2019megatron) 26.04 20.98
@@
Comparison with Other Layer Normalization. In addition, we conducted comparisons using LLaMA-130M to evaluate LayerNorm Scaling against recently proposed normalization methods, including Admin (liu2020understanding), Sandwich-LN (ding2021cogview), Group-LN (wu2018group; ma2024megalodon), and Mix-LN (li2024mix). Table 7 shows that Admin and Group-LN degrade performance. Sandwich-LN slightly outperforms Pre-LN. Both Mix-LN and LayerNorm Scaling improve over Pre-LN by good margins. However, Mix-LN fails to reduce perplexity under 26, falling short of LayerNorm Scaling and suffers from instability in large-scale scenarios as shown in Table 1.
@@ 7 Related Work
Ineffectiveness of Deeper Layers in Transformers. The ineffectiveness of deep layers in LLMs has been previously reported. yin2023outlier found that deeper layers of LLMs can tolerate significantly higher levels of pruning compared to shallower layers, achieving high sparsity. Similarly, gromov2024unreasonable and men2024shortgpt demonstrated that removing early layers causes a dramatic decline in model performance, whereas removing deep layers does not. lad2024remarkable showed that the middle and deep layers of GPT-2 and Pythia exhibit remarkable robustness to perturbations such as layer swapping and layer dropping. Recently, li2024owlore highlighted that early layers contain more outliers and are therefore more critical for fine-tuning. While these studies effectively highlight the limitations of deep layers in LLMs, they stop short of identifying the root cause of this issue or proposing viable solutions to address it.
@@
Layer Normalization in Language Models. LN (ba2016layer) was initially applied after the residual connection in the original Transformer (vaswani2017attention), which is known as Post-LN. Later on, Pre-LN (baevski2018adaptive; dai2019transformer; nguyen2019transformers) dominated LLMs, due to its compelling performance and stability (brown2020language; touvron2023llama; jiang2023mistral; bi2024deepseek). Prior works have studied the effect of Pre-LN and Post-LN. xiong2020layer proves that Post-LN tends to have larger gradients near the output layer, which necessitates smaller learning rates to stabilize training, whereas Pre-LN scales down gradients with the depth of the model, working better for deep Transformers. wang2019learning empirically confirmed that Pre-LN facilitates stacking more layers and Post-LN suffers from gradient vanishing. The idea of connecting multiple layers was proposed in previous works (bapna2018training; dou2018exploiting; wang2019learning). Admin introduces additional parameters to control residual dependencies, stabilizing Post-LN. DeepNorm (wang2024deepnet) enables stacking 1000-layer Transformers by upscaling the residual connection before applying LN. Additionally, ding2021cogview proposed Sandwich LayerNorm, normalizing both the input and output of each transformer sub-layer. takase2022b2t introduced B2T to bypass all LN except the final one in each layer. li2024mix recently combines Post-LN and Pre-LN to enhance the middle layers. zhu2025transformersnormalization introduces Dynamic Tanh (DyT) as a normalization-free alternative in Transformers, delivering comparable performance. zhuo2025hybridnormstableefficienttransformer proposes HybridNorm, a hybrid normalization scheme combining QKV normalization with Post-Norm FFN to stabilize training in deep transformers. de2020batchnormalizationbiasesresidual also states that normalized residual blocks in deep networks are close to the identity function and proposes SkipInit to remove normalization by introducing a learnable scalar multiplier on the residual branch initialized to $1/\sqrt{L}$. Our experiments suggest that SkipInit’s learnable parameter does not improve performance and sometimes harms training.
@@ Appendix A Proofs of the Theorems of curse of depth
Given Equation (2) from takase2023spike, we have:
@@ A.1.2 Variance of the Feed-Forward Network
Using the result obtained by wang2024deepnet, we get:
@@
This provides a tighter lower bound for $\sigma^{2}_{x_{\ell}}$} compared to the upper bound of Equation (33). Since we know the upper bound of variance grows exponentially, the lower bound must be sub-exponential. Therefore, for $\sigma^{2}_{x_{\ell}}$$\ell$, we must have:
@@
∎Therefore, the increasing lower bound for $\sigma^{2}_{x_{\ell}}$} must grow faster than a linear function. So, the increase of variance is sub-exponential. A large increase in such bound will lead to gradient spikes, which can connect to previous studies in huang2025stable; huang2025spam.
@@
$\times\Big(1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\Big(\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big)\Big).$
@@ A.2.1 Proof of Lemma 36
Our derivation follows results in takase2023spike, specifically Equation (7), which provides an upper bound on the norm of $\frac{\partial y_{\ell}}{\partial x_{1}}$ as:
@@
$\text{FFN:}\left\|\frac{\partial y_{\ell}}{\partial x^{\prime}_{\ell}}\right\|_{2}\quad\text{Attention:}\left\|\frac{\partial x^{\prime}_{\ell}}{\partial x_{\ell}}\right\|_{2}.$ (38)
@@
Let $\sigma_{w1_{\ell}}$ and $\sigma_{w2_{\ell}}$ be the standard deviations of $W^{1}_{\ell}$ and $W^{2}_{\ell}$, respectively. From Assumption 3.1, the spectral norms of $W^{1}_{\ell}$ and $W^{2}_{\ell}$ are given by their standard deviations and dimensions (vershynin2018), so wo have:
@@
From vershynin2018, we know that:
@@
$\times\Big(1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\Big(\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big)\Big).$
@@ A.2.2 Analysis of the Upper Bound
As discussed in takase2023spike, $\sigma$ should be sufficiently small, and the standard deviation, $\sigma_{x^{\prime}_{\ell}}$ or $\sigma_{x_{\ell}}$ should satisfy the condition $\sigma^{2}\ll\sigma_{x^{\prime}_{\ell}}$ to maintain the lazy training scheme. Thus, we obtain the following bound for the product over $\ell$ from 1 to $L$:
@@
Equation (36) can be expressed in a traditional product form Whittaker_Watson_1996 for $\sigma_{x_{\ell}}$:
@@
To establish a lower bound for $\sigma^{2}_{x_{\ell}}$}, we exponentiate both sides. Setting $\sigma^{2}_{x_{\ell}}$$\ell$, we must have:
@@
Here $\epsilon$ is a small constant with $1/2\leq\epsilon<1$. Therefore, the increasing upper bound of $\sigma^{2}_{x_{\ell}}$} is slower than the $\ell^{3}$ function, leading to: